# RAGServe: Fast Quality-Aware RAG Systems with Configuration Adaptation

### v1

@@ 10 Conclusion
This paper introduces RAGServe, the first system that focuses on optimizing the tradeoffs between response delay and generation quality in RAG, by by jointly scheduling RAG queries and adapting key configurations on a per-query basis. Evaluation on four datasets shows that RAGServe outperforms the state-of-the-art, reducing generation latency by 1.642.54×1.64-2.54$\times$1.64 - 2.54 × without compromising response quality.
@@ Appendix A Appendix
We use a very simple prompt to provide the metadata to RAGServe’ LLM profiler. We don’t perform any prompt tuning or optimizations.
@@
4 metadata = "ThedatasetconsistsofmultiplechunksofinformationfromFortune500companiesonfinancialreportsfromeveryquarterof2023."
@@
The chunk_size is chosen based on guidelines RAG literature for different types of RAG tasks [24, 49]. We don’t tune this knob as it is fixed when the database is created.

### v3

@@ 10. Conclusion
This paper introduces METIS, the first system that focuses on optimizing the tradeoffs between response delay and generation quality in RAG, by by jointly scheduling RAG queries and adapting key configurations on a per-query basis. Evaluation on four datasets shows that METIS outperforms the state-of-the-art, reducing generation latency by $1.64-2.54\times$ without compromising response quality.
Acknowledgment
We thank all the anonymous reviewers and our shepherd Oana Balmau, for their insightful feedback and suggestions. The project is funded by NSF CNS-2146496, CNS-2131826, CNS-2313190, CNS-1901466, UChicago CERES Center and a Google Faculty Research Award.
@@ Appendix A Appendix
The appendix has not been not peer-reviewed.
A.1. Prompt and input syntax for METIS’ LLM profiler
We use a simple prompt to provide the metadata to METIS’ LLM profiler. We don’t perform any prompt tuning or optimizations for this work as the goal of the prompt was only to get binary decisions from the natural language properties of the query.
@@ A.1. Prompt and input syntax for METIS’ LLM profiler
4 metadata = "ThedatasetconsistsofmultiplechunksofinformationfromFortune500companiesonfinancialreportsfromeveryquarterof2023.Thechunksizeis1024tokens."
5
@@
The chunk_size is chosen based on guidelines RAG literature for different types of RAG tasks [24, 60]. We don’t tune this knob as it is fixed when the database is created. Finally in this work, we don’t tune the metadata for the dataset, we use the existing summaries.
Today, RAG QA datasets already have summaries present along with the queries and contexts. In future work , it will be interesting to study how to effectively construct such a metadata for newer datasets. One possible solution could be an LLM summarizer on a set of values from the dataset which opens up further avenues to perform joint scheduling and configuration tuning.
A.2. Changing the embedding algorithm for the vector database in METIS
METIS picks a state-of-art retrieval algorithm Cohere-embed-v3.0 [4]. Using two other popular retrieval algorithms All-mpnet-base-v2 [72] and text-embedding-3-large-256 [18], the F1-score change remained within 1%. The delay has no measurable difference as the retrieval is $100\times$ faster than LLM synthesis [6].